OpenStudy (nikki_skellington):

Wan has 22 bulbs of the same shape and size in a box. The colors of and amounts of the bulbs are shown below: 6 blue bulbs 9 red bulbs 7 orange bulbs Without looking in the box, Wan takes out a bulb at random. He then replaces the bulb and takes out another bulb from the box. What is the probability that Wan takes out an orange bulb in both draws? 7 / 22 * 7 / 22 = 49 / 484 7 / 22 * 6 / 21 = 42 / 462 7 /22 + 6 / 21 = 279 / 462 7 / 22 + 7 / 22 = 308 / 484

OpenStudy (nikki_skellington):

i think this one is C (7 /22 + 6 / 21 = 279 / 462) am i right?

jimthompson5910 (jim_thompson5910):

A = probability of orange = (# of orange)/(# total) = 7/22 probability of 2 oranges = (A)*(A) = (7/22)*(7/22) this only works if you put the bulb back

OpenStudy (nikki_skellington):

ok thank you, again lol

jimthompson5910 (jim_thompson5910):

np